Using the sample space for rolling two dice, construct a probability distribution for the random variable X representing the sum of the numbers that appear.

1
Expert's answer
2022-03-13T15:54:13-0400

 

The sample space "\varOmega" is

 "\left( 1,1 \right) ,\left( 1,2 \right) ,\left( 1,3 \right) ,\left( 1,4 \right) ,\left( 1,5 \right) ,\left( 1,6 \right) \\\left( 2,1 \right) ,\left( 2,2 \right) ,\left( 2,3 \right) ,\left( 2,4 \right) ,\left( 2,5 \right) ,\left( 2,6 \right) \\\left( 3,1 \right) ,\left( 3,2 \right) ,\left( 3,3 \right) ,\left( 3,4 \right) ,\left( 3,5 \right) ,\left( 3,6 \right) \\\left( 4,1 \right) ,\left( 4,2 \right) ,\left( 4,3 \right) ,\left( 4,4 \right) ,\left( 4,5 \right) ,\left( 4,6 \right) \\\left( 5,1 \right) ,\left( 5,2 \right) ,\left( 5,3 \right) ,\left( 5,4 \right) ,\left( 5,5 \right) ,\left( 5,6 \right) \\\left( 6,1 \right) ,\left( 6,2 \right) ,\left( 6,3 \right) ,\left( 6,4 \right) ,\left( 6,5 \right) ,\left( 6,6 \right)"

The sum for each "\omega \in \varOmega" is respectively

"2,3,4,5,6,7\\3,4,5,6,7,8\\4,5,6,7,8,9\\5,6,7,8,9,10\\6,7,8,9,10,11\\7,8,9,10,11,12"

Hence for X - sum of rolls:

"P\left( X=2 \right) =\frac{1}{36}\\P\left( X=3 \right) =\frac{2}{36}=\frac{1}{18}\\P\left( X=4 \right) =\frac{3}{36}=\frac{1}{12}\\P\left( X=5 \right) =\frac{4}{36}=\frac{1}{9}\\P\left( X=6 \right) =\frac{5}{36}\\P\left( X=7 \right) =\frac{6}{36}=\frac{1}{6}\\P\left( X=8 \right) =\frac{5}{36}\\P\left( X=9 \right) =\frac{4}{36}=\frac{1}{9}\\P\left( X=10 \right) =\frac{3}{36}=\frac{1}{12}\\P\left( X=11 \right) =\frac{2}{36}=\frac{1}{18}\\P\left( X=12 \right) =\frac{1}{36}"